# A lower bound on the number of edges in DP-critical graphs

### v1

@@ 4.3 Subgraphs induced by low vertices in $G$
Suppose that ${\mathcal{B}}_{0}$ contains $C=\{v_{1},\ldots,v_{b}\}$ such that $B:=G[C]$ is a ${(k-2)}$-regular block. Let $M$ be the set of $(k-2)$-vertices in $C$. For all $v_{i}\in C-M$, let $v^{\prime}_{i}$ be the unique neighbor of $v_{i}$ outside of $C$. Since $G$ is $h$-minimal, $G-C$ has an $(H,L)$-coloring $f$. We write $H_{f}$ for the set of colors in $H$ above $B$ that are adjacent to $f$. For each $v_{i}\in C$, we write $L_{f}(v_{i})=L(v_{i})\setminus V(H_{f})$. Since $|L_{f}(v)|=d_{B}(v)$ for each $v\in V(B)$ and $B$ is not $(H[L_{f}(C)],L_{f})$-colorable, Theorem 1.7 implies that $B$ is a multiple of either a cycle of a complete graph.
If $B$ is isomorphic to $K_{k-1}$, then by Lemma 3.1, $H[L_{f}(C)]$ consists of $k-2$ copies of the graph $K_{k-1}$. In particular, $H[L_{f}(C)]$ is $(k-2)$-regular. By Lemmas 4.17, $M$ is nonempty, so $H_{f}$ is a complete graph on at most $k-2$ vertices; in particular, $H_{f}$ is not $(k-2)$-regular.
If $B$ is not isomorphic to $K_{k-1}$, then by Lemma 3.1, every component of $H[L_{f}(C)]$ is a $(k-2)$-regular $q$-blowup of a cycle or a complete graph for some $q$. More precisely, writing $t$ for the regularity of $\widetilde{B}$, $q=\frac{k-2}{t}$. Therefore, $H_{f}$ is a subgraph of $\widetilde{B}$, and hence each vertex of $H_{f}$ has degree at most $t<k-2$. In particular, $H_{f}$ is not $(k-2)$-regular.
As $H_{f}$ is a component of $H[L_{f}(C)]$, which is not $(k-2)$-regular,
$\forall v_{i}\in C-M$, the color $f(v_{i}^{\prime})\in L(v^{\prime}_{i})$ is adjacent to the unique color $\alpha_{v_{i}}\in L(v_{i})\cap H_{f}$. (16)
@@
Fix a vertex $v_{i}\in C-M$, let $G_{i}=G-C$, and let $L_{i}$ differ from $L|_{V(G)-C}$ only in that $f(v_{i}^{\prime})$ is deleted from $L(v^{\prime}_{i})$. Correspondingly, $h_{i}$ differs from $h$ only in that $h_{i}(v^{\prime}_{i})=h(v^{\prime}_{i})-1$. By (16), $G_{i}$ is not $(H[L_{i}(V(G_{i}))],L_{i})$-colorable. So, there is an $h_{i}$-minimal subgraph $G^{\prime}_{i}$ of $G_{i}$, and it contains $v^{\prime}_{i}$. Since $h_{i}(v^{\prime}_{i})\leq k-2$, $G^{\prime}_{i}\neq K_{k-1}$, and $\rho_{G_{i},h_{i}}(V(G^{\prime}_{i}))\leq-2$. Then $\rho_{G,h}(V(G^{\prime}_{i}))\leq-2+(\lambda+2)=\lambda.$ So, by Lemma 4.5, there is a unique edge joining $V(G^{\prime}_{i})$ with $V(G)\setminus V(G^{\prime}_{i})$. Since $C\cap V(G^{\prime}_{i})=\emptyset$, this edge is $v_{i}v^{\prime}_{i}$. Hence each edge connecting $C$ with $V(G)-C$ is a cut edge. As $B^{*}$ is $2$-edge-connected, it follows that $x_{0}^{*}\in C$, contradicting Lemma 4.16. ∎
5 Discharging
@@
  1. 1.
    For each $v\in V(G)$, we give initial charge $\rho_{h}(v)$ to $v$. For each pair $uv$, where $u,v\in V(G)$ are joined by $s\geq 1$ edges, we give initial charge $-s(2\lambda+1)+1$ to the pair $uv$. For each non-adjacent pair $u,v\in V(G)$, the initial charge of $uv$ is 00.
  2. 2.
@@
Now, we investigate the final charges of the vertices in $G$. For each $v\in V(G)$, we write $ch^{*}(v)$ for the final charge of $v$. We observe that the total charge in $G$ is $\rho_{h}(G)$. Additionally, the final charge of each vertex pair is 00, so $\rho_{h}(G)=\sum_{v\in V(G)}ch^{*}(v)$. Finally, if $S^{*}_{0}\neq V(G)$, then the total charge in $G[\overline{S^{*}_{0}}]$ is at most $\rho_{G,h}(\overline{S^{*}_{0}})-\lambda\leq 0$ by Lemma 4.8. Therefore,
$\rho_{h}(G)=\sum_{v\in V(G)}ch^{*}(v)\leq\sum_{v\in S^{*}_{0}}ch^{*}(v).$ (17)
@@
For each $v\in S^{*}_{0}$, if $d(v)=h(v)=j$ for some fixed $j\in\{3,\ldots,k-1\}$, then we consider two cases:
@@
    $ch^{*}(v)\leq(j\lambda-1)-j\lambda-\frac{d(v)-\widetilde{d}(v)}{2}-\alpha(d_{B^{*}}(v)-d_{{\mathcal{B}}_{0}}(v))=-1-\frac{d(v)-\widetilde{d}(v)}{2}-\alpha(d_{B^{*}}(v)-d_{{\mathcal{B}}_{0}}(v)).$
@@
    If $j=k-1$, then
    $ch^{*}(v)$$\leq$ $(k-1)\lambda+1-(k-1)\lambda-\frac{d(v)-\widetilde{d}(v)}{2}-\alpha(d_{B^{*}}(v)-d_{{\mathcal{B}}_{0}}(v))$
    $=$ $1-\frac{d(v)-\widetilde{d}(v)}{2}-\alpha d_{B^{*}}(v)+\alpha d_{{\mathcal{B}}_{0}}(v).$
We claim that ${\mathcal{B}}_{0}$ is nonempty. Indeed, suppose that ${\mathcal{B}}_{0}$ is empty. As $d(v)\geq 1+h(v)$ for each $v\in S_{0}^{*}$ we have the following inequality instead of (N1)–(N3):
@@
Then, by (17), $\rho_{h}(G)<-2|S^{*}_{0}|\leq-2$, so $G$ is not a counterexample. Thus, we assume that ${\mathcal{B}}_{0}\neq\emptyset$, or equivalently, that $S_{0}^{*}$ contains at least one low vertex.
@@
If $B$ is a component of ${\mathcal{B}}_{0}$, then $\sum_{v\in V(B)}ch^{*}(v)<-1$.
@@
We note that $d_{B^{*}}(v)=h(v)$ for $v\in V(B)\setminus\{x^{*}_{0}\}$, and $d_{B^{*}}(x^{*}_{0})=h(x^{*}_{0})-1$ whenever $x^{*}_{0}\in V({\mathcal{B}}_{0})$. Hence, by (L1)–(L2),
$\sum_{v\in V(B)}ch^{*}(v)\leq|V_{k-1}(B)|-|V_{k-1}^{-}(B)|-\sum_{v\in V(B)}\left(\frac{d(v)-\widetilde{d}(v)}{2}\right)-\alpha\sum_{v\in V(B)}\left(h(v)-d_{{\mathcal{B}}_{0}}(v)\right)+\alpha$
$=|V_{k-1}(B)|-|V_{k-1}^{-}(B)|-m(B)-\alpha\sigma_{h}(B)+\alpha=-\Phi_{k}(B)+\alpha.$
Note that the $\alpha$ term accounts for the possibility that $x^{*}_{0}\in V(B)$, in which case $-d_{B^{*}}(x^{*}_{0})=-h(x^{*}_{0})+1$.
By Corollary 4.12, Condition (i) of Lemma 3.2 holds for $T=B$. Condition (ii) holds because the vertices in $B$ are low. By Lemma 4.18, $B$ has no $(k-2)$-regular block. If $B$ is $(k-1)$-regular, then as $B$ is low and $G$ is connected, $G=B$, which contradicts Lemma 4.13. Therefore, Condition (iii) holds. So, by Lemma 3.2, $\sum_{v\in V(B)}ch^{*}(v)\leq-\Phi_{k}(B)+\alpha<-1$. ∎
Now, by (N1)–(N3), the vertices of $S_{0}^{*}\setminus{\mathcal{B}}_{0}$ have total nonpositive charge. Therefore,
$\rho_{h}(G)=\sum_{v\in V(G)}ch^{*}(v)\leq\sum_{v\in S^{*}_{0}}ch^{*}(v)\leq\sum_{v\in{\mathcal{B}}_{0}}ch^{*}(v).$
As ${\mathcal{B}}_{0}\neq\emptyset$, Lemma 5.1 implies that $\sum_{v\in{\mathcal{B}}_{0}}ch^{*}(v)<-1$. Therefore, $\rho_{h}(G)<-1$, and as $\rho_{h}(G)$ is integral, $\rho_{h}(G)\leq-2$. This completes the proof.
6 Bound for list coloring
@@
Theorem 6.1.
Let $k\geq 5$, and let $G$ be a multigraph with a function $h:V(G)\to\{0,\dots,k-1\}$. Suppose that $G$ has no exceptional subgraph. If $(H,L)$ is an $h$-cover of $G$ and $G$ is $(H,L)$-minimal, then $\rho_{h}(G)\leq-2$.
@@
Suppose for the sake of contradiction that $\rho_{h}(G)\geq-1$. As $G$ is $(H,L)$-minimal, no vertex $v\in V(G)$ satisfies $h(v)=0$. By Theorem 2.1, $G$ is not $h$-minimal. Since $G$ is not DP $h$-colorable, $G$ has a proper subgraph $G_{0}$ that is $h$-minimal. Since $G$ has no exceptional subgraph, Theorem 2.1 implies that $\rho_{G,h}(V(G_{0}))\leq-2$. We let $G_{1}\supseteq G_{0}$ be a maximal subgraph of $G$ for which $\rho_{G,h}(V(G_{1}))\leq-2$, and we observe that $V(G_{1})\subsetneq V(G)$. By the $(H,L)$-minimality of $G$, $G_{1}$ has an $(H,L)$-coloring $f$.
@@
$-1\leq\rho_{G,h}(V(G_{1})\cup U)\leq-2-2j\lambda+\rho_{G,h}(U).$
Rearranging, $\rho_{G,h}(U)\geq 1+2j\lambda$, which implies that $\rho_{G,h^{\prime}}(U)\geq 1+j(\lambda-2)$. Thus, every subgraph $G^{\prime\prime}$ of $G^{\prime}$ satisfies $\rho_{h^{\prime}}(G^{\prime\prime})\geq 1$, and hence by Theorem 2.1, $G^{\prime}$ has no $h^{\prime}$-critical subgraph. Therefore, $G^{\prime}$ is DP $h^{\prime}$-colorable, and thus $f$ can be extended to an $(H,L)$-coloring of $G$. This contradicts the original assumption that $G$ is $(H,L)$-minimal, completing the proof. ∎
Corollary 6.2.
@@
Now, we construct a $(k-1)$-cover $(H,L)$ of $G$ for which the $L^{\prime}$-coloring and the $(H,L)$-coloring problems on $G$ are equivalent. As $G$ has no $L^{\prime}$-coloring and is list $k$-critical, it follows that $G$ is $(H,L)$-minimal. Therefore, by Theorem 6.1, $\rho_{h}(G)=(k\lambda+1)n-2\lambda m\leq-2$. Rearranging, $m\geq(k-1+\frac{1}{\lambda})\frac{n}{2}+\frac{1}{\lambda}$. ∎

### v2

@@ 4.3. Subgraphs induced by low vertices in $G$
Suppose that ${\Lambda}_{0}$ contains $C=\{v_{1},v_{2},\ldots,v_{b}\}$ such that $B:=G[C]$ is a (k2){(k-2)}-regular block. Let $M$ be the set of (k2)(k-2)-vertices in $C$. For all $v_{i}\in C-M$, let $v^{\prime}_{i}$ be the unique neighbor of $v_{i}$ outside of $C$. Since $G$ is $h$-minimal, GCG-C has an $(H,L)$-coloring $\varphi$. We write $H_{\varphi}$ for the set of colors in $H$ above $B$ that are adjacent to $\varphi$. For each $v_{i}\in C$, we write $L_{\varphi}(v_{i})=L(v_{i})\setminus V(H_{\varphi})$. Since $|L_{\varphi}(v)|=d_{B}(v)$ for each $v\in V(B)$ and $B$ is not $(H[L_{\varphi}(C)],L_{\varphi})$-colorable, Theorem 1.5 implies that $B$ is a multiple of either a cycle of a complete graph.
If $B$ is isomorphic to $K_{k-1}$, then by Lemma 3.1, $H[L_{\varphi}(C)]$ consists of k2k-2 copies of the graph $K_{k-1}$. In particular, $H[L_{\varphi}(C)]$ is (k2)(k-2)-regular. By Lemma 4.17, $M$ is nonempty, so $H_{\varphi}$ is a complete graph on at most k2k-2 vertices; in particular, $H_{\varphi}$ is not (k2)(k-2)-regular.
If $B$ is not isomorphic to $K_{k-1}$, then by Lemma 3.1, every component of $H[L_{\varphi}(C)]$ is a (k2)(k-2)-regular $q$-blowup of a cycle or a complete graph for some $q$. More precisely, writing $t$ for the regularity of $\widetilde{B}$, $q=\frac{k-2}{t}$. Therefore, $H_{\varphi}$ is a subgraph of $\widetilde{B}$, and hence each vertex of $H_{\varphi}$ has degree at most t<k2t<k-2. In particular, $H_{\varphi}$ is not (k2)(k-2)-regular.
As $H_{\varphi}$ is a component of $H[L_{\varphi}(C)]$, which is not (k2)(k-2)-regular,
(17) $\forall v_{i}\in C-M$, the color $\varphi(v_{i}^{\prime})\in L(v^{\prime}_{i})$ is adjacent to the unique color $\alpha_{v_{i}}\in L(v_{i})\cap H_{\varphi}$.
@@
Fix a vertex $v_{i}\in C-M$, let $G_{i}=G-C$, and let $L_{i}$ differ from $L|_{V(G)-C}$ only in that $\varphi(v_{i}^{\prime})$ is deleted from $L(v^{\prime}_{i})$. Correspondingly, $h_{i}$ differs from $h$ only in that $h_{i}(v^{\prime}_{i})=h(v^{\prime}_{i})-1$. By (17), $G_{i}$ is not $(H[L_{i}(V(G_{i}))],L_{i})$-colorable. So, there is an $h_{i}$-minimal subgraph $G^{\prime}_{i}$ of $G_{i}$, and it contains $v^{\prime}_{i}$. Since $h_{i}(v^{\prime}_{i})\leq k-2$, $G^{\prime}_{i}\neq K_{k-1}$, and $\rho_{G_{i},h_{i}}(V(G^{\prime}_{i}))\leq-2$. Then $\rho_{G,h}(V(G^{\prime}_{i}))\leq-2+(\lambda+2)=\lambda.$ So, by Lemma 4.5, there is a unique edge joining $V(G^{\prime}_{i})$ with $V(G)\setminus V(G^{\prime}_{i})$. Since $C\cap V(G^{\prime}_{i})=\emptyset$, this edge is $v_{i}v^{\prime}_{i}$. Hence each edge connecting $C$ with V(G)CV(G)-C is a cut edge. As $B^{*}$ is $2$-edge-connected, it follows that $x_{0}^{*}\in C$, contradicting Lemma 4.16. ∎
5. Discharging
@@
  1. (1)
    For each $v\in V(G)$, we give initial charge $\rho_{h}(v)$ to $v$. For each pair $uv$, where $u,v\in V(G)$ are joined by $s\geq 1$ edges, we give initial charge $s(2\lambda+1)+1$ to the pair $uv$. For each pair $uv$ of non-adjacent $u,v\in V(G)$, the initial charge of $uv$ is 0.
  2. (2)
@@
Now, we consider the final charges of the vertices in $G$. For each $v\in V(G)$, we write $ch^{*}(v)$ for the final charge of $v$. We observe that the total charge in $G$ is $\rho_{h}(G)$. Additionally, the final charge of each vertex pair is 0, so $\rho_{h}(G)=\sum_{v\in V(G)}ch^{*}(v)$. Finally, if $S^{*}_{0}\neq V(G)$, then the total charge in $G[\overline{S^{*}_{0}}]$ is at most $\rho_{G,h}(\overline{S^{*}_{0}})-\lambda\leq 0$ by Lemma 4.8. Therefore,
(18) $\rho_{h}(G)=\sum_{v\in V(G)}ch^{*}(v)\leq\sum_{v\in S^{*}_{0}}ch^{*}(v).$
@@
For each $v\in S^{*}_{0}$, if $d(v)=h(v)=j$ for some fixed $j\in[3,\ldots,k-1]$, then we consider two cases:
@@
    $ch^{*}(v)\leq(j\lambda-1)-j\lambda-\frac{d(v)-\widetilde{d}(v)}{2}-\alpha(d_{B^{*}}(v)-d_{{\Lambda}_{0}}(v))=-1-\frac{d(v)-\widetilde{d}(v)}{2}-\alpha(d_{B^{*}}(v)-d_{{\Lambda}_{0}}(v)).$
@@
    If j=k1j=k-1, then
    $ch^{*}(v)$ $\leq$ $(k-1)\lambda+1-(k-1)\lambda-\frac{d(v)-\widetilde{d}(v)}{2}-\alpha(d_{B^{*}}(v)-d_{{\Lambda}_{0}}(v))$
    $=$ $1-\frac{d(v)-\widetilde{d}(v)}{2}-\alpha d_{B^{*}}(v)+\alpha d_{{\Lambda}_{0}}(v).$
We claim that ${\Lambda}_{0}$ is nonempty. Indeed, suppose that ${\Lambda}_{0}$ is empty. As $d(v)\geq 1+h(v)$ for each $v\in S_{0}^{*}$ we have the following inequality instead of (N1)–(N3):
@@
Then, by (18), $\rho_{h}(G)<-2|S^{*}_{0}|\leq-2$, so $G$ is not a counterexample. Thus, we assume that ${\Lambda}_{0}\neq\emptyset$, or equivalently, that $S_{0}^{*}$ contains at least one low vertex.
@@
If $B$ is a component of ${\Lambda}_{0}$, then $\sum_{v\in V(B)}ch^{*}(v)<-1$.
@@
We note that $d_{B^{*}}(v)=h(v)$ for $v\in V(B)\setminus\{x^{*}_{0}\}$, and $d_{B^{*}}(x^{*}_{0})=h(x^{*}_{0})-1$ whenever $x^{*}_{0}\in V({\Lambda}_{0})$. Hence, by (L1)–(L2),
$\sum_{v\in V(B)}ch^{*}(v)\leq|V_{k-1}(B)|-|V_{k-1}^{-}(B)|-\sum_{v\in V(B)}\left(\frac{d(v)-\widetilde{d}(v)}{2}\right)-\alpha\sum_{v\in V(B)}\left(h(v)-d_{{\Lambda}_{0}}(v)\right)+\alpha$
$|V_{k-1}(B)|-|V_{k-1}^{-}(B)|-m(B)-\alpha\sigma_{h}(B)+\alpha=-\Phi_{k}(B)+\alpha.$
Note that the $\alpha$ term accounts for the possibility that $x^{*}_{0}\in V(B)$, in which case $d_{B^{*}}(x^{*}_{0})=-h(x^{*}_{0})+1$.
By Corollary 4.12, Condition (i) of Lemma 3.2 holds for $T=B$. Condition (ii) holds because the vertices in $B$ are low. By Lemma 4.18, $B$ has no (k2)(k-2)-regular block. If $B$ is (k1)(k-1)-regular, then as $B$ is low and $G$ is connected, $G=B$, which contradicts Lemma 4.13. Therefore, Condition (iii) holds. So, by Lemma 3.2, $\sum_{v\in V(B)}ch^{*}(v)\leq-\Phi_{k}(B)+\alpha<-1$. ∎
Now, by (N1)–(N3), the vertices of $S_{0}^{*}\setminus{\Lambda}_{0}$ have total nonpositive charge. Therefore,
$\rho_{h}(G)=\sum_{v\in V(G)}ch^{*}(v)\leq\sum_{v\in S^{*}_{0}}ch^{*}(v)\leq\sum_{v\in{\Lambda}_{0}}ch^{*}(v).$
As ${\Lambda}_{0}\neq\emptyset$, Lemma 5.1 implies that $\sum_{v\in{\Lambda}_{0}}ch^{*}(v)<-1$. Therefore, $\rho_{h}(G)<-1$, and as $\rho_{h}(G)$ is integral, $\rho_{h}(G)\leq-2$. This completes the proof.
6. Bound for list coloring
@@
Lemma 6.1.
Let $k\geq 5$, and let $G$ be a multigraph with a function $h:V(G)\to[0,k-1]$. Suppose that $G$ has no exceptional subgraph. If $(H,L)$ is an $h$-cover of $G$ and $G$ is $(H,L)$-minimal, then $\rho_{h}(G)\leq-2$.
@@
Suppose for the sake of contradiction that $\rho_{h}(G)\geq-1$. As $G$ is $(H,L)$-minimal, no vertex $v\in V(G)$ satisfies $h(v)=0$. By Theorem 2.1, $G$ is not $h$-minimal. Since $G$ is not DP $h$-colorable, $G$ has a proper subgraph $G_{0}$ that is $h$-minimal. Since $G$ has no exceptional subgraph, Theorem 2.1 implies that $\rho_{G,h}(V(G_{0}))\leq-2$. We let $G_{1}\supseteq G_{0}$ be a maximal subgraph of $G$ for which $\rho_{G,h}(V(G_{1}))\leq-2$, and we observe that $V(G_{1})\subsetneq V(G)$. By the $(H,L)$-minimality of $G$, $G_{1}$ has an $(H,L)$-coloring $\varphi$.
@@
$1\leq\rho_{G,h}(V(G_{1})\cup U)\leq-2-2j\lambda+\rho_{G,h}(U).$
Rearranging, $\rho_{G,h}(U)\geq 1+2j\lambda$, which implies that $\rho_{G,h^{\prime}}(U)\geq 1+j(\lambda-2)$. Thus, every subgraph $G^{\prime\prime}$ of $G^{\prime}$ satisfies $\rho_{h^{\prime}}(G^{\prime\prime})\geq 1$, and hence by Theorem 2.1, $G^{\prime}$ has no $h^{\prime}$-critical subgraph. Therefore, $G^{\prime}$ is DP $h^{\prime}$-colorable, and thus $\varphi$ can be extended to an $(H,L)$-coloring of $G$. This contradicts the original assumption that $G$ is $(H,L)$-minimal, completing the proof. ∎
Now, with Lemma 6.1, we can prove Theorem 1.7.
Theorem 1.7.
@@
Now, we construct a (k1)(k-1)-cover $(H,L)$ of $G$ for which the $L^{\prime}$-coloring and the $(H,L)$-coloring problems on $G$ are equivalent. As $G$ has no $L^{\prime}$-coloring and is list $k$-critical, it follows that $G$ is $(H,L)$-minimal. Therefore, by Lemma 6.1, $\rho_{h}(G)=(k\lambda+1)n-2\lambda m\leq-2$. Rearranging, $m\geq(k-1+\frac{1}{\lambda})\frac{n}{2}+\frac{1}{\lambda}$. ∎
Concluding remarks. We do not think that our bounds are sharp. On the other hand, it seems that $f_{\mathrm{DP}}(n,k)<f(n,k)$ for many values of $n$ and $k$. An interesting problem is to find the asymptotics of $\frac{f_{\mathrm{DP}}(n,k)}{n}$ for a fixed $k$ and $n\to\infty$.